an iClicker operates at a frequency of approximately 900MHz (900*10^6) what is the approximate wavelength of the electromagnetic wave produced by the iClicker
Answer #1

We know

Wavelength = c/frequency

C is speed of light = 3*10^8

Wavelength = 3*10^8/900*10^6

Frequency = 1/3 =. 33 second^(-1)